it Is often suggested that the Bank of Canada tries to achieve zero inflation. If we assume that velocity is constant, does this zero inflation goal require that the rate of money growth equal zero? If yes, why? If no, then what should the rate of money growth equal?
1
Expert's answer
2020-04-01T09:25:42-0400

The quantitative theory of money says that there is a direct correlation between the amount of money in the economy and the price level of goods and services sold.

"P=\frac{M\times V}{Q}"

p - price level

M - money supply

V - money circulation rate, how many times a year money is turned around in the economy

Q - real GNP


Let V = 2, Q = 500, then we get:

"0=\frac{M\times2}{500}"

M=0